OpenStudy (anonymous):

For which \(a\) does the following series converge? \[\sum_{n=2}^{\infty}\frac{1}{n(\ln n)^a}\]

OpenStudy (anonymous):

For which \(a\) does the following series converge? \[\sum_{n=2}^{\infty}\frac{1}{n(\ln n)^a}\]

OpenStudy (amistre64):

if i recall correctly; fractions add up and converge? id say for all a > 0 but i cant be sure if

OpenStudy (anonymous):

consider the harmonic series \[\frac{1}{n^1}\] does not converge

OpenStudy (amistre64):

yeah, im still going over series and convergence divergence material in the book :) itll sink in one of these days